OpenStudy (anonymous):

Find the center and the radius of the equation below. I don't think my answer is correct.. i got (1,-2) for the center and 1=radius. can someone check for me please? the equation is below

OpenStudy (anonymous):

\[4x^2+4y^2-8x+24y+4=0\]

OpenStudy (isaiah.feynman):

group the x's and y's together, and complete the square separately.

OpenStudy (anonymous):

that's what i did :) but my circle doesn't look.... like a circle. to find more points, don't i just move 1 unite up, down, left, right because of the radius??

OpenStudy (anonymous):

Let us first simplify by dividing all terms by 4,. That way you will have x^2 and y^2 with a coefficient of 1

OpenStudy (anonymous):

oooh forgot that! thanks @Florcelda !

OpenStudy (anonymous):

show me the new expression, and we will go from there, to group x and y terms

OpenStudy (anonymous):

new expression is 4(x^2+y^2-2x+6y+1)

OpenStudy (anonymous):

=zero

OpenStudy (anonymous):

we can divide both sides by 4 giving x^2 + y^2 - 2x + 6y + 1 = 0

OpenStudy (anonymous):

so is my standard form (x-1)^2+(y+3)^2=8??

OpenStudy (anonymous):

you can now transpose 1 to the right side and group our x's and y's. Leave a space for the linear terms. We will complete the squares

OpenStudy (anonymous):

sorry, i'm moving a little ahead of you :)

OpenStudy (anonymous):

= 9

OpenStudy (anonymous):

did you found your mistake in getting 8?

OpenStudy (anonymous):

oh i see, i forgot to add the +1 from the xterm, thank you! so this makes my circle (1,-3) and radius 3?

OpenStudy (anonymous):

Perfect!

OpenStudy (anonymous):

thanks! so in order to find other points to make sure.. i just use the radius ( go up, down, left and right 3units from the center?)

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

My pleasure